OpenStudy (anonymous):

I want to post a 3*3 matrix...with elements a11=1 a12=0 a13=0 a21=1 a22=0 a23=1 a31=0 a32=1 a33=0...what is the syntax

OpenStudy (unklerhaukus):

\[\text{\[ \left( \begin{array}\ 1 & 0 &0 \\ 1 &0& 1\\ 0& 1 & 0 \end{array}\right)\]}\]

OpenStudy (anonymous):

@UnkleRhaukus how did you type this reply without making it appear as a matrix...

OpenStudy (unklerhaukus):

\[\text{\[\text{\[\left(\begin{array}\1&0&0\\1&0&1\\0&1&0\end{array}\right)\]}\]}\]